# New Tool to Detect Inhomogeneous Chiral Symmetry Breaking

### v2

@@ Abstract
In this letter, we discuss a novel method to search for inhomogeneous chiral symmetry breaking in theories with fermions. The prime application we have in mind is QCD, but the method is also applicable for other theories, including solid state applications. It is based on an extension of the chiral susceptibility to inhomogeneous phases and it works as a stability analysis. Our method allows us to determine when homogeneous solutions have the tendency to create spatially modulated condensates. As proof of principle, we apply this technique to a rainbow-ladder QCD model and confirm the conjectured shape of the instability region published in Ref. [1].
@@ I Introduction
Chiral Symmetry Breaking (SB) is one of the most consequential phenomena associated with strong interactions. With massless quarks, the QCD Lagrangian is manifestly chirally symmetric, however the ground state breaks chiral symmetry. For two massless quark flavours, at high temperatures the system undergoes a second-order phase transition to the symmetric phase, which becomes a cross-over at finite quark masses [2, 3]. What happens to chiral symmetry at large chemical potentials, however, is less clear. One possibility is that the system undergoes a transition to a phase where chiral symmetry is broken inhomogeneously, that is, the order parameter is a periodic function of space. These phases have been seen in models of strong interactions [4, 5, 6, 7]. If real, these phases might occur in a region of the phase diagram that is connected to the physics of neutron stars and neutron-star-mergers as well as with low-energy heavy-ion collisions. In fact, it has already been argued [8, 9, 10, 11] that moat regimes (a broader but correlated phenomenon to inhomogeneous phases, see also [12, 13]) would leave experimental signatures in scattering experiments.
@@
From a theoretical perspective, it is highly non-trivial to obtain solid information on the existence of inhomogeneous phases at large chemical potential. Lattice calculations of QCD are notoriously difficult if not prohibited in this region because of the sign problem [14]. Therefore, some attention has been paid to simpler (and partly lower-dimensional) models, such as Gross-Neveu (GN) or Nambu-Jona-Lasinio (NJL) type models [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], which are accessible to both, continuum and lattice methods. Although much has been learned from the models, in the end one needs to study the problem within QCD.
@@
In general, two strategies have been employed so far. Using a specific ansatz for the shape of the inhomogeneous phase, one can calculate the free energy in this configuration and compare with the homogeneous one. One of the most commonly used ansatzes, e.g., is the so-called Chiral Density Wave (CDW) [5]. Here, the condensates oscillate with spatial variable
@@
where is the third Pauli matrix in flavour space and is the wave vector of the modulation. In the NJL-model the amplitude of this modulation depends on the coupling constant to the scalar and pseudo-scalar sectors [5]. Alternatively, one can perform what is called a stability analysis which probes the stability of the homogeneous phase with respect to very small inhomogeneous perturbations. This technique has the benefit that it is agnostic with respect to the shape of the inhomogeneous phase. On the other hand, since it is based on small perturbations, it could miss instabilities with respect to large fluctuations. Nevertheless, these two techniques typically agree on second-order phase boundaries.
@@
Both, the ansatz and the stability analysis techniques face severe additional difficulties in QCD compared to their application in simple models. Due to the non-locality of the self-energy the expectation values of quark bilinears are functions of two momentum variables in an inhomogeneous phase. To our knowledge, Ref. [30] is the only attempt so far to deal with these difficulties in a functional approach to QCD via Dyson-Schwinger Equations (DSEs). Using an ansatz for the quark propagator which yields condensates like Eq. (1), they show that spatially inhomogeneous solutions to their set of (truncated) DSEs appear at high chemical potentials.
@@
The stability analysis framework has only recently been generalised from model applications to QCD [31, 1]. Although agnostic in general, our specific formulation of the framework made necessary by the added layer of complexity of QCD relied on the specific shape of the perturbations (what we called the “test-function”). Furthermore, we were only able to apply the analysis to a very limited region of the phase diagram, namely the left spinodal line. The details are explained in Refs. [31, 1], here we just note that the shape of our test-functions has constraints, which we can only exactly enforce in this region. These shortcomings notwithstanding, the general methodology of stability analysis is still highly promising and its first application to a simple truncation of QCD [1] shows the tendency to create inhomogeneous condensates.
@@
In this letter, we overcome these shortcomings by introducing a novel technique to perform a stability analysis. This new method is applicable to any theory, including simple models, arbitrary truncations of QCD but also effective theories in solid state physics. The technique does not depend on any test-functions, and it is not restricted to any region of the phase diagram. It is based on a well-known method to identify the second-order homogeneous phase boundaries, namely calculating the chiral susceptibility (see e.g. Refs. [32, 33]). However, by performing a chiral rotation to the quark fields, we are able to compute an inhomogeneous chiral susceptibility and determine when the symmetric solution is unstable with respect to inhomogeneous chiral symmetry breaking.
The outline of the Letter is as follows. In the next section we will introduce this idea in the context of NJL. We show how it reproduces previous results. We then move to generalise it to QCD and show first proof-of-principle results in a simple truncation. Discussions and conclusions follow.
@@ II Inhomogeneous Chiral Susceptibility in NJL
Let us illustrate the idea in the NJL model. Take the following standard Lagrangian (note that we will use the same Euclidian-space conventions as in Ref. [31] and as for the numerical values for the couplings, cutoff, and regularisation scheme, we will take the setup of Ref. [4])
We will work in the chiral limit of zero bare mass . In mean-field NJL, the inverse dressed fermion propagator
where is the dynamically generated mass, which depends on the chiral condensate, and is an order parameter for SB. One way to resolve the phase diagram of the theory is to calculate the derivative of this order parameter with respect to the bare quark mass. For the symmetric phase, this is simply
where we denote the combination of the Matsubara sum of the frequencies and 3-momentum integration as . The susceptibility is negative where the symmetric phase is unstable with respect to small homogeneous breakings of chiral symmetry, and it will be positive otherwise. Close to a second-order chiral transition, the sign change of exactly finds the phase transition. Around a first-order transition, however, there is a region where both the restored and the broken solution are (meta-)stable. This is known as the spinodal region and its edges are known as the spinodal lines. In Figure 1, we show how the sign of neatly resolves the second-order transition and left-spinodal lines in NJL, i.e., the regions where the restored solution changes from unstable to stable.
Take now the following (chiral-rotated) modified quark fields [34]
The NJL-Lagrangian, Eq. (2), then transforms to
For non-zero , assuming a homogeneous condensate and , it is easy to show that we obtain a modulated just like Eq. (1),
Note that, when we omit the spatial arguments in the condensate, we mean that it is constant in space.
Under this transformation, the inverse fermion propagator for non-zero becomes
This again can be inverted and an inhomogeneous chiral susceptibility can be calculated to be the following
This expression is general, however, in practice we take the wave-vector taken to be fixed in the direction, i.e., . In Fig. 2 we show how this resolves exactly the phase boundary of the inhomogeneous phase as well: By varying we manage to find where the symmetric solution is unstable with respect to inhomogeneous SB, in regions where it is not unstable with respect to homogeneous SB. The convolution of the red regions in Fig. 2, that is, regions of negative for some value of , reproduces exactly the unstable region in Ref. [4].
Let us now proceed to extending this idea to truncations of QCD.
@@ III QCD
As it was the case for the NJL example, first we need to find the chiral symmetric solution, in order to calculate its inhomogeneous chiral susceptibility. We take a rainbow-ladder model of QCD defined by the following quark Dyson-Schwinger Equation
@@
Here, the bare line denotes the (inverse) bare propagator, whereas the line with the black dot represents the dressed (inverse) propagator including quantum fluctuations,
where . The self-interaction diagram contains a dressed gluon propagator (wiggly line) and two bare quark-gluon vertices. The effect of dressing the gluon propagator with momentum and one of the quark gluon vertices, , is modelled by two analytic functions and
Here, factors of the QCD coupling and renormalisation functions of the vertex and the quark have been absorbed into the effective couplings and . The transverse () and longitudinal () projectors are given by
As an example for such an effective running coupling we take the Qin-Chang model [35] with a thermal mass for the gluon calculated via hard-thermal and hard-dense loop techniques
This model has been applied to aspects of the QCD phase diagram before, see e.g. the review Ref. [36] for a summary.
As a first step in the analysis, we solve the DSE for the homogeneous restored phase, for which is exactly zero, all over the phase diagram. This is done in the standard formulation of QCD without the chiral rotation Eq. (5). We then force the appearance of a modulation by performing the rotation, which adds a term to the inverse propagator, i.e., in the QCD Lagrangian
where here refers to the covariant derivative. In the propagator, this extension induces several new terms which can be worked out exactly. Their details do not matter here. Each of these new terms is accompanied by its own dressing function. In an ansatz-based approach, one would have to go through the complicated and extremely costly exercise to solve the resulting set of DSEs self-consistently to get the stable CDW solution.
Is is, however, much simpler and numerically cheaper to determine the inhomogeneous chiral susceptibility in this framework. A trivial extrapolation of the calculation above will show that the inhomogeneous chiral susceptibility of the chiral solution is the following
@@
Again, a non-trivial modulated chiral condensate is indicated if this function goes negative at finite . Comparing Eq. (17) with the corresponding one for the NJL-model, Eq. (9), we find the slight complication that the susceptibility now depends on the quark momentum and appears on both sides, and the equation has to be solved self-consistently. Note, however, that the quark dressing functions and on the right hand side are the ones calculated without the extra term present in the theory. This is a subtle and important point. We are not seeking for a self-consistent solution of the DSE with the extra term present. Instead, the question we want answered is: once we add , will this induce the formation of a Dirac scalar term to the quark propagator? If so, then the susceptibility diverges at the second-order phase boundary and an inhomogeneous phase will be formed.
In Fig. 3 we show the line where the susceptibility diverges for different values of . This analysis not only confirms our earlier findings from Ref. [1] but also significantly advances our results by moving beyond test-function-based stability analyses. Here, we conclusively show that the chiral solution is unstable to inhomogeneous perturbations, with a proto-Lifschitz point 111A proto-Lifschitz point is a term we define in Ref. [1] and it simply refers to the point where the stability boundary of the symmetric phase w.r.t. inhomogeneous fluctuations meets the left spinodal of the homogeneous first-order transition appearing near MeV. As seen in Fig. 3, we can definitively delineate the region bounded by the spinodal line on one side and the inhomogeneous convolution (black dotted) line on the other as a region where the chiral solution really is unstable against crystallisation. Importantly, we establish that for temperatures as low as 70 MeV, the instability region does not extend into the first-order transition. Since in this region the restored solution is not energetically favoured over the homogeneous broken solution, we identify this as a feature of a disfavored phase rather than an indicator of an emergent inhomogeneous phase. Note, however, that with this method we now can perform the stability analysis where the homogeneous restored solution is preferred over the homogeneous broken one, and, there, we identify that it is stable w.r.t. small inhomgeneities.
Exploring lower temperatures may reveal points where the instability line crosses the first-order phase transition. However, our current methodology—though substantially more precise and conclusive than prior approaches—becomes increasingly resource-intensive at lower temperatures since it requires solving the homogeneous DSE with large numerical precision. Thus, rather than extending this model analysis further, we prioritize more realistic truncations as a direction for future work, where our methodology can provide more rigorous insights. This work establishes a solid foundation and methodology to probe these complexities in the future, with higher-fidelity models.
@@ IV Conclusions
In this work, we proposed a new method to analyse the stability of a chirally symmetric phase against the formation of an inhomogeneous condensate via chiral symmetry breaking. The method is based on an extension of the concept of chiral susceptibility toward probing the reaction of the theory to the presence of spatially modulated quark fields. As a proof of principle, we showed that our method quantitatively reproduces previous NJL-model results [4] on the presence of inhomogeneous phases and we confirmed previous expectations for instability in a rainbow-ladder model of QCD [1].
We would like to stress again, that this methodology to study inhomogeneous chiral symmetry breaking is very general. It does not depend on a test-function and it is equally applicable to any truncation of QCD, with no computational costs coming from complex truncations. Since it is not necessary to know the full momentum dependence of , we expect this should also be easily calculable with the functional renormalisation group.
Some studies suggest that additional quantum fluctuations could destabilise inhomogeneous condensates (see Ref. [38]). In that case, we expect the homogeneous solution to the DSE to capture the energetically preferred, stable configuration. Other studies indicate that the boundary between inhomogeneous and restored phases may be first-order (see Ref. [39]), limiting our ability to precisely determine the transition line. Nevertheless, our approach should still identify the corresponding spinodal.
@@
One should, however, keep in mind that a stability analysis in general does not say much about the fully self-consistent solution of the system. It just shows that there is the tendency to create a mass term (break chiral symmetry) once we introduce a modulation, in this instance a CDW-like condensation. The actual shape of the modulation would eventually to be determined by a self-consistent solution of the DSEs for all dressing functions of the quark. This is a tremendous effort, but one that should be undertaken once our method does show instabilities in more advanced truncations to QCD than used in this work.

### v3

@@ Abstract
In this letter, we discuss a novel method to search for inhomogeneous chiral symmetry breaking in theories with fermions. The prime application we have in mind is QCD, but the method is also applicable for other theories, including solid state applications. It is based on an extension of the chiral susceptibility to inhomogeneous phases and it works as a stability analysis. Our method allows us to determine when homogeneous solutions have the tendency to create spatially modulated condensates. As proof of principle, we apply this technique to a rainbow-ladder QCD model and find that its phase diagram contains an inhomogeneous region.
@@ I Introduction
Chiral Symmetry Breaking (SB) is one of the most consequential phenomena associated with strong interactions. With massless quarks, the QCD Lagrangian is manifestly chirally symmetric, however the ground state breaks this symmetry. For two massless quark flavours, at high temperatures the system undergoes a second-order phase transition to the symmetric phase, which becomes a cross-over at finite quark masses [1, 2]. What happens to chiral symmetry at large chemical potentials, however, is less clear. One possibility is that the system undergoes a transition to a phase where chiral symmetry is broken inhomogeneously, that is, the order parameter is a periodic function of space. These phases have been seen in models of strong interactions [3, 4, 5, 6]. If real, these phases might occur in a region of the phase diagram that is connected to the physics of neutron stars and neutron-star-mergers as well as with low-energy heavy-ion collisions. In fact, it has already been argued [7, 8, 9, 10] that moat regimes (a broader but correlated phenomenon to inhomogeneous phases, see also [11, 12]) would leave experimental signatures in scattering experiments.
@@
From a theoretical perspective, it is highly non-trivial to obtain solid information on the existence of inhomogeneous phases at large chemical potential. Lattice calculations of QCD are notoriously difficult if not prohibited in this region because of the sign problem [13]. Therefore, some attention has been paid to simpler (and partly lower-dimensional) models, such as Gross-Neveu or Nambu-Jona-Lasinio (NJL) type models [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], which are accessible to both, continuum and lattice methods.
@@
In general, two strategies have been employed so far. Using a specific ansatz for the shape of the inhomogeneous phase, one can calculate the free energy in this configuration and compare with the homogeneous one. One of the most commonly used ansatzes is the so-called Chiral Density Wave (CDW) [4]111The CDW is closely related to the (Quarkyonic) Chiral Spiral in dimensions [29, 30]. Other commonly used ansatz functions are the Real-Kink-Crystal [20, 3], the Twisted Kink Crystal [31, 32], and similar hybrid condensates [33].. Here, the condensates oscillate with spatial variable
@@
where is the third Pauli matrix in flavour space and is the wave vector of the modulation. Alternatively, one can perform what is called a stability analysis which probes the stability of the homogeneous phase with respect to very small inhomogeneous perturbations. This technique has the benefit that it is agnostic with respect to the shape of the inhomogeneous phase. On the other hand, since it is based on small perturbations, it could miss instabilities with respect to large fluctuations. Nevertheless, these two techniques typically agree on second-order phase boundaries.
@@
Although much has been learned from models, in the end one needs to study the problem within QCD. Both, the ansatz and the stability analysis techniques face severe additional difficulties in QCD compared to their application in simple models. Due to the non-locality of the self-energy the expectation values of quark bilinears are functions of two momentum variables in an inhomogeneous phase. To our knowledge, Ref. [34] is the only attempt so far to deal with these difficulties in a functional approach to QCD via Dyson-Schwinger Equations (DSEs). Using an ansatz for the quark propagator which yields condensates like Eq. (1), they show that spatially inhomogeneous solutions to their set of (truncated) DSEs appear at high chemical potentials.
@@
The stability analysis framework has only recently been generalised from model applications to QCD [35, 36]. Although agnostic in general, the QCD formulation of the framework relied on the specific shape of the perturbations (what is called the “test-function”). Furthermore, so far, it has been possible to be applied to a very limited region of the phase diagram, namely the left spinodal line. The details are explained in Refs. [35, 36], here we just note that the shape of the test-functions has constraints, which one can exactly enforce only in this region. Thus, although this version of the stability analysis is very promising and has led to first evidence of a inhomogeneous phase in a simple truncation of QCD [36], it suffers somewhat from these limitations.
@@
In this letter, we overcome these shortcomings by introducing an alternative and novel technique to perform a stability analysis. The technique does not depend on any test-functions, and it is not restricted to any region of the phase diagram. It is based on a well-known method to identify the second-order homogeneous phase boundaries, namely calculating the chiral susceptibility (see e.g. Refs. [37, 38]). However, by performing a space-dependent chiral rotation of the quark fields, we are able to compute an inhomogeneous chiral susceptibility and determine when the symmetric solution is unstable with respect to inhomogeneous chiral symmetry breaking.
@@ II Inhomogeneous Chiral Susceptibility in NJL
Before we generalize it to QCD, let us illustrate the idea in the NJL model, defined by the Lagrangian222We will use the same Euclidian-space conventions as in Ref. [35] and as for the numerical values for the couplings, cutoff, and regularisation scheme, we will take the setup of Ref. [3].
We will work in the chiral limit of zero bare mass . In mean-field, the inverse dressed fermion propagator reads
where is the dynamically generated mass, which depends on the chiral condensate and is an order parameter for SB. One way to resolve the phase diagram of the theory is to calculate the derivative of this order parameter with respect to the bare quark mass. For the symmetric phase () and colors, this is simply
where we denote the combination of the Matsubara sum of the frequencies and 3-momentum integration as . The susceptibility is negative where the symmetric phase is unstable with respect to small homogeneous breakings of chiral symmetry, and it is positive otherwise. In a second-order chiral transition, changes its sign and diverges on the phase boundary. Around a first-order transition, however, there is a region where both the restored and the broken solution are (meta-)stable. This is known as the spinodal region and its edges are known as the spinodal lines. In Fig. 1, we show how the sign of neatly resolves the second-order transition and left-spinodal lines in NJL, i.e., the regions where the restored solution changes from unstable to (meta-)stable.
Take now the following chiral-rotated quark fields [39]
The NJL-Lagrangian, Eq. (2) with , then transforms to
For non-zero , assuming a homogeneous condensate and , it is easy to show that one obtains a modulated just like Eq. (1),
Under this transformation, the inverse fermion propagator becomes
It can still be inverted and an inhomogeneous chiral susceptibility in the symmetric phase can be calculated333For this we temporarily re-introduce a small bare quark mass to the gap equation after the chiral rotation and send it to zero after taking the derivative. according to
This expression is general, however, in practice we take the wave-vector taken to be fixed in the direction, i.e., . In Fig. 2 we show how this resolves exactly the phase boundary of the inhomogeneous phase as well: By varying we manage to find where the symmetric solution is unstable with respect to inhomogeneous SB, in regions where it is not unstable with respect to homogeneous SB. The union of the red regions in Fig. 2, that is, regions of negative for some value of , reproduces exactly the unstable region in Ref. [3]. Note that the denominator of Eq. (9) is exactly the stability condition found in standard stability analysis, so this correspondence is exact.
Having verified that our new method works and reproduces known results within the NJL-model, let us now proceed extending this idea to QCD.
@@ III QCD
In QCD the setup is completely analogous. Under the chiral rotation, Eq. (5), the QCD Lagrangian
with covariant derivative and gluon field-strength tensor , transforms to
As in NJL, the procedure is to first find the chiral symmetric solution for the quark propagator of the non-rotated system, described by Eq. (10). To this end we will use the DSE framework. Different approaches to obtain fully dressed n-point functions, as e.g. the functional renormalisation group are equally suitable. In this letter, in order to illustrate our new method we employ a simple rainbow-ladder truncation of the DSE for the quark propagator, given by
@@
Here, the bare line denotes the (inverse) bare quark propagator, whereas the line with the black dot represents the (inverse) dressed quark propagator including quantum fluctuations. It has the general structure
with two chiral-symmetry preserving terms parametrized by dressing functions and (where ) and a symmetry-breaking dressing function which vanishes in the symmetric phase. The self-interaction diagram contains two bare quark-gluon vertices. The dressing of one of these vertices has already been combined with the dressing of the gluon propagator (wiggly line) to an effective running coupling. For this coupling we will use the Qin-Chang model [40] in the same setup as detailed in [36]. This type of model has been applied to aspects of the QCD phase diagram before, see e.g. the review Ref. [41].
We now induce the inhomogeneous modulation by the chiral rotation. In an ansatz-based approach, one would have to go through the complicated and extremely costly exercise of solving the resulting set of DSEs self-consistently to get the stable CDW solution. On the contrary, it is much simpler and numerically cheaper to determine the inhomogeneous chiral susceptibility . Proceeding analogously to the NJL case yields the following expression (in the symmetric phase)
@@
where are the transverse and longitudinal components of the effective running coupling of the Qin-Chang model (see Ref. [36] for details). Again, a non-trivial modulated chiral condensate is indicated if goes negative at finite . Comparing Eq. (14) with the corresponding one for the NJL-model, Eq. (9), we find the slight complication that the susceptibility now depends on the quark momentum and appears on both sides such that the equation has to be solved self-consistently. Note, however, that the quark dressing functions and on the right hand side do not change under the chiral rotation in the symmetric phase and are therefore the same as those calculated without the extra term present in the theory. This is a subtle and important point. In this analysis, we do not need to solve the DSE with the extra term present. Instead, we study whether the addition of induces the formation of a Dirac scalar term to the quark propagator. If so, then the susceptibility diverges at the second-order phase boundary and an inhomogeneous phase will be formed.
In Fig. 3 we show the lines where the susceptibility diverges for different values of . We confirm our finding from Ref. [36] that near MeV there is a “proto-Lifschitz point” (pLP), i.e., a point where the stability boundary of the symmetric phase w.r.t. inhomogeneous fluctuations meets the left spinodal of the homogeneous first-order transition. Below this point there is a region where the symmetric phase is unstable against developing inhomogeneous structures. However, this instability is without consequences as long as it occurs to the left of the first-order homogeneous phase boundary (blue solid line), since in this region the symmetric phase is also disfavored against the chiral broken phase. In Ref. [36] the technical complexity of the applied test-function method prevented us from moving away very far from the left spinodal, and therefore we could only speculate about the existence of an inhomogeneous phase at higher chemical potential. Our new method, on the other hand, can be applied at arbitrarily high chemical potential. Varying the value of we find a triple-point (3P) at MeV where the instability line of the symmetric phase crosses the homogeneous first-order phase boundary. In particular we establish that, while for temperatures above 60 MeV, the instability region does not extend beyond the homogeneous first-order transition, at lower temperatures it does. Hence, in this regime the symmetric phase is unstable against inhomogeneous fluctuations in a region where it is favoured over the homogeneous chiral broken solution. This is an unambiguous proof that, in this truncation, a crystalline phase would surface.
Although this is very exciting, the main point of this study is to demonstrate that we now have a methodology where this analysis can be performed. Our ultimate goal is to prioritize more realistic truncations as a direction for future work, where our methodology can provide more rigorous insights.
@@ IV Conclusions
In this work, we proposed a new method to analyse the stability of a chirally symmetric phase against the formation of an inhomogeneous condensate via chiral symmetry breaking. The method is based on an extension of the concept of chiral susceptibility toward probing the reaction of the theory to the presence of spatially modulated quark fields. As a proof of principle, we showed that our method exactly reproduces previous NJL-model results [3] on the presence of inhomogeneous phases, we confirmed previous results for the presence of a proto-Lifschitz Point in a rainbow-ladder model of QCD [36], and we now also find a true instability of the symmetric phase in a regime where it is favored over the homogeneous broken phase.
However, the main point of this letter is not this specific result but rather that the proposed method is indeed suited to perform such analyses. In fact, with this methodology we were able to obtain substantially more conclusive results than with the test-function based stability analysis applied in Ref. [36]. The latter requires numerically searching for a saddle point in the function’s parameter space, which can be very expensive. Moreover, choosing the parameter space too small bears the risk of wrong results. For that reason this analysis is so far only feasible close to the left spinodal line [36] where the sufficiency of the parameter space can be ensured by analysing the homogeneous transition. Our new method requires no external parameters, and the only real numerically costly operation is that of solving the homogeneous DSE.
@@
The only restriction of the method is that, as it is based on a chiral rotation, it primarily signals instabilities toward CDW-like condensates, see Eq. (7). However, as known from model studies, even though the CDW is typically not the most favored inhomogeneous phase, in the chiral limit its second-order phase boundary to the symmetric phase coincides with those for other modulations. In this context one should keep in mind that a stability analysis in general does not say much about the fully self-consistent solution of the system. Here it just shows that there is the tendency to break chiral symmetry once we introduce a modulation, in this instance a CDW-like condensation. The actual shape of the modulation would eventually to be determined by a self-consistent solution of the DSEs for all dressing functions of the quark. This is a tremendous effort, but one that should be undertaken once our method does show instabilities in more advanced truncations to QCD than used in this work.
Some more care is required to generalise the method to be applicable away from the chiral limit. Then the chiral rotation (5) is not a symmetry transformation of the QCD Lagrangian but generates non-standard mass terms that need to be dealt with. This is work in progress.
On the other hand, the new method can easily be generalized to any theory, including simple models, but also effective theories in solid state physics, to study the inhomogeneous breaking of any exact symmetry of that theory. For instance, in the QCD context it could also be used to study inhomogeneous color superconductivity [42]. Also, it is equally applicable to any truncation of QCD, with no computational costs in addition to solving the homogeneous DSE. Since it is not necessary to know the full momentum dependence of , we expect this should also be calculable with the functional renormalisation group.
Some studies suggest that additional quantum fluctuations could destabilise inhomogeneous condensates [43, 44, 45, 30]. In that case, we expect the homogeneous solution to the DSE to capture the energetically preferred, stable configuration. Other studies indicate that the boundary between inhomogeneous and restored phases may be first order [46], limiting our ability to precisely determine the transition line. Nevertheless, our approach should still identify the corresponding spinodal.